# Imaging the positron annihilation line with 20-year INTEGRAL/SPI observations

### v2

@@ 3.1.3 Expected count calculation
The total expected counts consist of contributions from both the 511 keV sky map and the background components. Combining these components, the expected counts can be calculated as:
$\begin{split}\epsilon_{p,d,e}&=\epsilon^{\mathrm{source}}_{p,d,e}+\epsilon^{\mathrm{bkg}}_{p,d,e}\\ &=\sum_{j}R_{p,d,e,j}\lambda_{j}\\ &+b^{\mathrm{line}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{line}}_{p,d,e}+b^{\mathrm{cont}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{cont}}_{p,d,e}+\sum_{m}B^{\mathrm{source},m}_{p,d,e}\leavevmode\nobreak\ .\end{split}$ (5)
@@ 3.2 Image deconvolution
$\log L=\sum_{p,d,e}\left(D_{p,d,e}\log\epsilon_{p,d,e}-\epsilon_{p,d,e}\right)\leavevmode\nobreak\ .$ (6)
@@
The RL algorithm iteratively updates the sky map based on the comparison between the observed and expected counts. In each iteration, the update of the sky map is calculated as:
@@
$\lambda^{\mathrm{new}}_{j}$ $=\lambda^{\mathrm{old}}_{j}+\alpha\left[\omega_{j}\delta\lambda_{j}\right]_{\mathrm{smooth}}\leavevmode\nobreak\ ,$ (8)
where $\epsilon_{p,d,e}$ is the expected count calculated by Equation 5; $\alpha$ is a parameter to accelerate the RL algorithm. Here, the weight factor $\omega_{j}$ is introduced to enhance the reconstruction efficiency by providing stronger feedback to regions with longer exposure:
$\omega_{j}\equiv\left(\frac{\sum_{p,d,e}R_{p,d,e,j}}{\mathrm{max}\left(\sum_{p,d,e}R_{p,d,e,j}\right)}\right)^{l}\leavevmode\nobreak\ .$ (9)
The parameter $l$ controls the strength of this weighting. While the previous studies (Knödlseder et al. 2005; Siegert et al. 2020) adopted $l=0.5$, we explore various values to determine the optimal choice (see Section 3.3). Also, following Siegert et al. (2020), we adopt the operator $\cdot]_{\mathrm{smooth}}$ representing Gaussian smoothing applied to the weighted difference map. This smoothing suppresses high-frequency noise in the image, enhancing stability in the image reconstruction. However, a too-large smoothing kernel can lead to losing information about high-frequency structures in the image. We optimize the FWHM of this Gaussian kernel $\theta_{\mathrm{smooth}}$ with the data, which will be described in the next subsection.
@@
$b^{\mathrm{line/cont,new}}_{\mathrm{rev}}=\dfrac{b^{\mathrm{line/cont,old}}_{\mathrm{rev}}}{\displaystyle\sum_{p\in\sigma(\mathrm{rev}),d,e}B^{\mathrm{line/cont}}_{p,d,e}}\displaystyle\sum_{p\in\sigma(\mathrm{rev}),d,e}\dfrac{D_{p,d,e}}{\epsilon_{p,d,e}}B^{\mathrm{line/cont}}_{p,d,e}\leavevmode\nobreak\ ,$ (10)
where $\sigma(\mathrm{rev})$ denotes the set of pointings in a given orbital revolution. To assess the impact of this background optimization, we also perform reconstructions without updating these normalization factors, which were already optimized during the background modeling and data screening phase. This alternative approach is used to evaluate our systematic uncertainty, as explained in the following subsection.
The iteration process is terminated when the improvement in the log-likelihood ($\Delta L=\log L^{\mathrm{new}}-\log L^{\mathrm{old}}$) becomes smaller than a threshold. Here we set it to $10^{-2}$. It is known that running iterations until the likelihood converges can cause artifacts to appear in the reconstructed image. Previous studies often stopped the iterations early to focus only on global structures. However, using the 20-year dataset, we aim to search for new features in addition to well-known large-scale structures. Therefore, we continue the iterations until reaching reasonable convergence in the likelihood. To address the concern about the artifacts, we employ the bootstrap analysis to evaluate statistical fluctuations, allowing us to assess the uncertainties and significance of features of interest. Its details are presented in Section 3.4.
@@ 3.3 Parameter optimization and systematic error estimation
To find the optimal parameters, we performed the image reconstructions with various parameter combinations and identified one that maximizes the log-likelihood of the reconstructed image. Regarding the smoothing kernel size, we observe that $\theta_{\mathrm{smooth}}$ of 2 degrees yields the maximum log-likelihood, which is close to SPI’s angular resolution of 2.7 degrees. The weighting power $l$ shows optimal values between 0.5 and 1.0, with the maximum likelihood achieved at $l=0.75$. Note that the previous studies typically adopt $l=0.5$. Also, the reconstructed image is affected by $\lambda_{\mathrm{init}}$, which is the intensity of an initial flux map with units of cm-2 s-1 sr-1, as a starting point for the image reconstruction. A uniform intensity across the sky was assumed as the initial condition. The choice of $\lambda_{\mathrm{init}}$ influences the results, yielding different total fluxes of the reconstructed images. Based on comparison with the previous model-fitting studies (Skinner et al. 2015; Siegert et al. 2016), we adopted $\lambda_{\mathrm{init}}=5\times 10^{-5}$ cm-2 s-1 sr-1 as our optimal value because with this value the reconstructed total flux becomes consistent with the previous measurements best. The details of the parameter optimization are described in Appendix B.
The optimal parameter set determined from the above analysis is summarized in Table 2. Small deviations from the optimal parameters could potentially affect the reconstructed flux distribution, particularly in regions with lower exposure or faint signals. We performed additional image reconstructions to evaluate these systematic uncertainties by varying each parameter around its optimal value, as listed in Table 2. We also considered reconstructions with fixed background normalization, which has typically been employed in previous studies, as part of our systematic error evaluation. For each region of interest, we calculated fluxes from the reconstructed images with these different parameter sets and used their RMS to estimate their systematic uncertainty.
@@
We also evaluated systematic uncertainties from instrument response limitations, especially imperfections of the coded mask response function. We performed the image reconstruction with a modified response matrix of the coded mask aperture, where elements with effective areas of <<0.2 cm2 (corresponding to the edges and off-axis regions of the coded mask pattern) were set to zero. The resulting flux variations are $\lesssim$10% at most, especially in low-flux regions, and these uncertainties are also absorbed within other systematic errors. We note that photons penetrating the surrounding BGO shield could contribute a somewhat uniform background across all detectors, especially photons from outside $47^{\circ}\times 47\circ$, where the mask response is not covered in this analysis. This may potentially lead to flux misattribution. Quantitative evaluation of this effect would require detailed simulations of the entire payload geometry. While this is beyond the scope of this analysis, it should be noted that such systematic uncertainties are not accounted for in our current uncertainty estimation.
@@ 3.4 Bootstrap analysis
We perform a bootstrap analysis (Efron 1979) using two approaches to evaluate the statistical uncertainties of the fluxes of spatial features in the reconstructed image and their detection significance.
@@
Here, the background-induced fluctuations are estimated by using the same number of iterations as the reconstruction to be evaluated. Namely, for both approaches, we terminate the iterations at the same point, the 439th iteration, as used in the original data to ensure consistency in the analysis procedure. This is particularly important for the background-only bootstrap samples. If we adopt the convergence criterion described in Section 3.2, they often require more iterations to reconstruct an image with no signals, resulting in an underestimation of the background fluctuations in our case.
@@ 4.1 Global features
Figure 3: Reconstructed image of the 511 keV line emission from the 20-year INTEGRAL/SPI data. The image is shown in Galactic coordinates. Note that several dark spots appearing around the Galactic center, such as the one at $(l,b)\approx(353^{\circ},-4^{\circ})$, are artifacts often seen in the image deconvolution.
@@
The total flux integrated over the bulge region ($|l|<20^{\circ}$ and $|b|<20^{\circ}$) is measured to be $(1.36\pm_{\mathrm{stat}}0.04\pm_{\mathrm{sys}}0.05)\times 10^{-3}$ ph cm-2 s-1. The value after $\pm_{\mathrm{stat}}$ represents the statistical error calculated as the standard deviation of the flux distribution extracted from the bootstrapped samples, and the one after $\pm_{\mathrm{sys}}$ represents the systematic error derived from the results with different parameter sets (see Section 3.3). The measured flux shows agreement with the previous model-fitting results, as shown in Table 3. We note that our bulge flux measurement includes the disk emission overlapping with the bulge region, as we directly obtain the flux from the reconstructed image.
@@
Figure 4: Same as Figure 3 but with annotations showing the regions used for the flux measurements in Table 3.
@@ 4.2 Longitude and latitude profiles
To investigate the detailed spatial structure of the 511 keV emission, we examine flux profiles along both Galactic longitude and latitude. Figure 5 shows the longitude profile obtained by integrating the flux from $b$ = $20^{\circ}$ to $20^{\circ}$ in 4-degree longitude bins. While there is a slight flux enhancement in the negative longitude, this asymmetry is not statistically significant when considering both statistical and systematic uncertainties. To quantify this, we measured the fluxes in two specific regions, the eastern and western disk regions, as shown in Figure 4. The eastern disk ($l$ = $310^{\circ}-340^{\circ}$) shows the integrated flux of $(0.26\pm_{\mathrm{stat}}0.04\pm_{\mathrm{sys}}0.03)\times 10^{-3}$ ph cm-2 s-1, while the western disk ($l$ = $20^{\circ}-50^{\circ}$) yields $(0.19\pm_{\mathrm{stat}}0.03\pm_{\mathrm{sys}}0.03)\times 10^{-3}$ ph cm-2 s-1. The former is about 40% larger than the latter, but the statistical error is also $\sim$30%. Thus, the ratio between these regions is consistent with unity within current uncertainties. These results are also consistent with Skinner et al. (2015); Siegert et al. (2016), rather than the asymmetric profile claimed in early studies (Weidenspointner et al. 2008).
The latitude profile (Figure 6), derived by integrating the flux between $l=-20^{\circ}$ and $20^{\circ}$ in 4-degree latitude bins, shows potentially interesting features in the vertical structure of the 511 keV emission. The bulge emission exhibits different extensions at $|b|>10^{\circ}$. While the flux drops sharply at $b\approx-10^{\circ}$ in the northern region, the flux exhibits a plateau between $b=10^{\circ}$ and $\approx 18^{\circ}$. This corresponds to the more extended structure at ($l$, $b$) = ($10^{\circ}$$0^{\circ}$, $10^{\circ}$$20^{\circ}$) in the northeastern region of the bulge (see also Figure 3).
@@
Figure 6: Latitude profile of the 511 keV line emission obtained by integrating the flux from l=20l=-20 to +20+20 degrees in 4-degree latitude bins. The uncertainties are shown in the same manner as in Figure 5.
@@ 4.3.1 Search for the 511 keV from the star-forming regions
Figure 7: Same as Figure 3 but with a reduced maximum intensity of $1\times 10^{-3}$ ph cm-2 s-1 sr-1 to enhance the visibility of faint structures.
Figure 8: (Top) Same as Figure 7 but with annotations showing the regions used for flux measurements from individual sources (see Table 4). (Bottom) Same as the top panel but overlaid with contours of the sensitivity map shown in Figure 2.
Here, we investigate the 511 keV emission from individual regions, particularly the Cygnus region and the Scorpius–Centaurus OB association (see Figure 8). Including other star-forming regions, we evaluate the fluxes and their chance probabilities using the bootstrap samples with only the background.
@@
The Upper Scorpius region ($l=340^{\circ}$ to $360^{\circ}$, $b=25^{\circ}$ to $40^{\circ}$) is a part of the Scorpius–Centaurus association, the nearest OB association to the Sun. It shows a flux of $(1.4\pm_{\mathrm{stat}}0.8\pm_{\mathrm{sys}}0.5)\times 10^{-5}$ ph cm-2 s-1 and a chance probability of $8.0\times 10^{-2}$. This region corresponds to the chimney-like structure discussed in the previous subsection. Interestingly, 26Al maps measured by COMPTEL and INTEGRAL/SPI also show an enhanced emission from this region (Plüschke et al. 2001; Diehl et al. 2010). These findings suggest a possible spatial correlation between 26Al and positron annihilation. Considering that 26Al also emits a positron via its $\beta^{+}$ decay process, this correlation might shed light on the asymmetric vertical structure and the origin of the 511 keV emissions in this region with further confirmation. It will be discussed more in Section 5, also regarding the “OSSE 511 keV fountain”.
We also investigated the 511 keV emission from other prominent star-forming regions, the Vela, Orion-Eridanus, Perseus, and Orion regions. The Vela region ($l=260^{\circ}$ to $290^{\circ}$, $b=-15^{\circ}$ to $15^{\circ}$) shows a large chance probability of 19% with an upper limit of $8.3\times 10^{-5}$ ph cm-2 s-1. The Orion region ($l=190^{\circ}$ to $220^{\circ}$, $b=-50^{\circ}$ to $10^{\circ}$) similarly shows no significant excess when considered alone. When extending the region to include the Perseus OB association, the chance probability is somewhat improved to 10.1%. It should be noted that there could be possible contamination from the statistical fluctuation of the nearby bright source, the Crab. When we extracted the bright region ($l=170^{\circ}$ to $200^{\circ}$, $b=-30^{\circ}$ to $0^{\circ}$), its flux was derived to be $6.5\pm_{\mathrm{stat}}2.4\pm_{\mathrm{sys}}1.9$ ph cm-2 s-1 with an improved chance probability of $1.2\times 10^{-2}$. A faint signal may be seen at $l=120^{\circ}$ close to the Galactic plane in Figure 8, but its chance probability is about 60%, and we interpret it as a fluctuation in the reconstructed image.

### v3

@@ 3.1.3 Expected count calculation
The total expected counts consist of contributions from both the 511 keV sky map and the background components. Combining these components, the expected counts can be calculated as
$\begin{split}\epsilon_{p,d,e}&=\epsilon^{\mathrm{source}}_{p,d,e}+\epsilon^{\mathrm{bkg}}_{p,d,e}\\ &=\sum_{j}R_{p,d,e,j}\lambda_{j}\\ &+b^{\mathrm{line}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{line}}_{p,d,e}+b^{\mathrm{cont}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{cont}}_{p,d,e}+\sum_{m}B^{\mathrm{source},m}_{p,d,e}\penalty 10000\ .\end{split}$ (5)
@@ 3.2 Image deconvolution
$\log L=\sum_{p,d,e}\left(D_{p,d,e}\log\epsilon_{p,d,e}-\epsilon_{p,d,e}\right)\penalty 10000\ .$ (6)
@@
The RL algorithm iteratively updates the sky map based on the comparison between the observed and expected counts. In each iteration, the update of the sky map is calculated as
@@
$\lambda^{\mathrm{new}}_{j}$ $=\lambda^{\mathrm{old}}_{j}+\alpha\left[\omega_{j}\delta\lambda_{j}\right]_{\mathrm{smooth}}\penalty 10000\ ,$ (8)
where $\epsilon_{p,d,e}$ is the expected count calculated by Eq. 5; $\alpha$ is a parameter to accelerate the RL algorithm. Here, the weight factor $\omega_{j}$ is introduced to enhance the reconstruction efficiency by providing stronger feedback to regions with a longer exposure:
$\omega_{j}\equiv\left(\frac{\sum_{p,d,e}R_{p,d,e,j}}{\mathrm{max}\left(\sum_{p,d,e}R_{p,d,e,j}\right)}\right)^{l}\penalty 10000\ .$ (9)
The parameter $l$ controls the strength of this weighting. While the previous studies (Knödlseder et al. 2005; Siegert et al. 2020) adopted $l=0.5$, we explored various values to determine the optimal choice (see Sect. 3.3). Also, following Siegert et al. (2020), we adopt the operator $\cdot]_{\mathrm{smooth}}$ representing Gaussian smoothing applied to the weighted difference map. This smoothing suppresses high-frequency noise in the image, enhancing stability in the image reconstruction. However, a too-large smoothing kernel can lead to losing information about high-frequency structures in the image. We optimized the FWHM of this Gaussian kernel $\theta_{\mathrm{smooth}}$ with the data, which is described in the next subsection.
@@
$b^{\mathrm{line/cont,new}}_{\mathrm{rev}}=\dfrac{b^{\mathrm{line/cont,old}}_{\mathrm{rev}}}{\displaystyle\sum_{p\in\sigma(\mathrm{rev}),d,e}B^{\mathrm{line/cont}}_{p,d,e}}\displaystyle\sum_{p\in\sigma(\mathrm{rev}),d,e}\dfrac{D_{p,d,e}}{\epsilon_{p,d,e}}B^{\mathrm{line/cont}}_{p,d,e}\penalty 10000\ ,$ (10)
where $\sigma(\mathrm{rev})$ denotes the set of pointings in a given orbital revolution. To assess the impact of this background optimization, we also performed reconstructions without updating these normalization factors, which were already optimized during the background modeling and data screening phase. This alternative approach is used to evaluate our systematic uncertainty, as explained in the following subsection.
The iteration process is terminated when the improvement in the log-likelihood ($\Delta L=\log L^{\mathrm{new}}-\log L^{\mathrm{old}}$) becomes smaller than a threshold. Here we set it to $10^{-2}$. It is known that running iterations until the likelihood converges can cause artifacts to appear in the reconstructed image. Previous studies often stopped the iterations early to focus only on global structures. However, using the 20-year dataset, we aim to search for new features in addition to well-known large-scale structures. Therefore, we continued the iterations until reaching reasonable convergence in the likelihood. To address the concern about the artifacts, we employ the bootstrap analysis to evaluate statistical fluctuations, allowing us to assess the uncertainties and significance of features of interest. Its details are presented in Sect. 3.4.
@@ 3.3 Parameter optimization and systematic error estimation
To find the optimal parameters, we performed the image reconstructions with various parameter combinations and identified one that maximizes the log-likelihood of the reconstructed image. Regarding the smoothing kernel size, we observed that $\theta_{\mathrm{smooth}}$ of 2 degrees yielded the maximum log-likelihood, which is close to SPI’s angular resolution of 2.7 degrees. The weighting power $l$ shows optimal values between 0.5 and 1.0, with the maximum likelihood achieved at $l=0.75$. We note that the previous studies typically adopt $l=0.5$. Also, the reconstructed image is affected by $\lambda_{\mathrm{init}}$, which is the intensity of an initial flux map with units of cm-2 s-1 sr-1, as a starting point for the image reconstruction. A uniform intensity across the sky was assumed as the initial condition. The choice of $\lambda_{\mathrm{init}}$ influences the results, yielding different total fluxes of the reconstructed images. Based on comparison with the previous model-fitting studies (Skinner et al. 2015; Siegert et al. 2016), we adopted $\lambda_{\mathrm{init}}=5\times 10^{-5}$ cm-2 s-1 sr-1 as our optimal value because with this value the reconstructed total flux becomes consistent with the previous measurements best. The details of the parameter optimization are described in Appendix B.
The optimal parameter set determined from the above analysis is summarized in Table 2. Small deviations from the optimal parameters could potentially affect the reconstructed flux distribution, particularly in regions with lower exposure or faint signals. We performed additional image reconstructions to evaluate these systematic uncertainties by varying each parameter around its optimal value, as listed in Table 2. We also considered reconstructions with fixed background normalization, which has typically been employed in previous studies, as part of our systematic error evaluation. For each region of interest, we calculated fluxes from the reconstructed images with these different parameter sets and used their root mean square to estimate their systematic uncertainty.
@@
We also evaluated systematic uncertainties from instrument response limitations, especially imperfections of the coded mask response function. We performed the image reconstruction with a modified response matrix of the coded mask aperture, where elements with effective areas of <<0.2 cm2 (corresponding to the edges and off-axis regions of the coded mask pattern) were set to zero. The resulting flux variations are $\lesssim$10% at most, especially in low-flux regions, and these uncertainties are also absorbed within other systematic errors. We note that photons penetrating the surrounding BGO shield could contribute a somewhat uniform background across all detectors, especially photons from outside $47^{\circ}\times 47^{\circ}$, where the mask response is not covered in this analysis. This may potentially lead to flux misattribution. Quantitative evaluation of this effect would require detailed simulations of the entire payload geometry. While this is beyond the scope of this analysis, it should be noted that such systematic uncertainties are not accounted for in our current uncertainty estimation.
@@ 3.4 Bootstrap analysis
We performed a bootstrap analysis (Efron 1979) using two approaches to evaluate the statistical uncertainties of the fluxes of spatial features in the reconstructed image and their detection significance.
@@
Here, the background-induced fluctuations are estimated by using the same number of iterations as the reconstruction to be evaluated. Namely, for both approaches, we terminate the iterations at the same point, the 439th iteration, as used in the original data to ensure consistency in the analysis procedure. This is particularly important for the background-only bootstrap samples. If we adopt the convergence criterion described in Sect. 3.2, they often require more iterations to reconstruct an image with no signals, resulting in an underestimation of the background fluctuations in our case.
@@ 4.1 Global features
Figure 3: Reconstructed image of the 511 keV line emission from the 20-year INTEGRAL/SPI data. The image is shown in Galactic coordinates. We note that several dark spots appearing around the Galactic center, such as the one at $(l,b)\approx(353^{\circ},-4^{\circ})$, are artifacts often seen in the image deconvolution.
@@
The total flux integrated over the bulge region ($|l|<20^{\circ}$ and $|b|<20^{\circ}$) is measured to be $(1.36\pm_{\mathrm{stat}}0.04\pm_{\mathrm{sys}}0.05)\times 10^{-3}$ ph cm-2 s-1. The value after $\pm_{\mathrm{stat}}$ represents the statistical error calculated as the standard deviation of the flux distribution extracted from the bootstrapped samples, and the one after $\pm_{\mathrm{sys}}$ represents the systematic error derived from the results with different parameter sets (see Sect. 3.3). The measured flux shows agreement with the previous model-fitting results, as shown in Table 3. We note that our bulge flux measurement includes the disk emission overlapping with the bulge region, as we directly obtain the flux from the reconstructed image.
@@
Figure 4: Same as Fig. 3 but with annotations showing the regions used for the flux measurements in Table 3.
@@ 4.2 Longitude and latitude profiles
To investigate the detailed spatial structure of the 511 keV emission, we examine flux profiles along both Galactic longitude and latitude. Figure 5 shows the longitude profile obtained by integrating the flux from $b$ = $20^{\circ}$ to $20^{\circ}$ in 4-degree longitude bins. While there is a slight flux enhancement in the negative longitude, this asymmetry is not statistically significant when considering both statistical and systematic uncertainties. To quantify this, we measured the fluxes in two specific regions, the eastern and western disk regions, as shown in Fig. 4. The eastern disk ($l$ = $310^{\circ}-340^{\circ}$) shows the integrated flux of $(0.26\pm_{\mathrm{stat}}0.04\pm_{\mathrm{sys}}0.03)\times 10^{-3}$ ph cm-2 s-1, while the western disk ($l$ = $20^{\circ}-50^{\circ}$) yields $(0.19\pm_{\mathrm{stat}}0.03\pm_{\mathrm{sys}}0.03)\times 10^{-3}$ ph cm-2 s-1. The former is about 40% larger than the latter, but the statistical error is also $\sim$30%. Thus, the ratio between these regions is consistent with unity within current uncertainties. These results are also consistent with Skinner et al. (2015); Siegert et al. (2016), rather than the asymmetric profile claimed in early studies (Weidenspointner et al. 2008).
The latitude profile (Fig. 6), derived by integrating the flux between $l=-20^{\circ}$ and $20^{\circ}$ in 4-degree latitude bins, shows potentially interesting features in the vertical structure of the 511 keV emission. The bulge emission exhibits different extensions at $|b|>10^{\circ}$. While the flux drops sharply at $b\approx-10^{\circ}$ in the northern region, the flux exhibits a plateau between $b=10^{\circ}$ and $\approx 18^{\circ}$. This corresponds to the more extended structure at ($l$, $b$) = ($10^{\circ}$$0^{\circ}$, $10^{\circ}$$20^{\circ}$) in the northeastern region of the bulge (see also Fig. 3).
@@
Figure 6: Latitude profile of the 511 keV line emission obtained by integrating the flux from l=20l=-20 to +20+20 degrees in 4-degree latitude bins. The uncertainties are shown in the same manner as in Fig. 5.
@@ 4.3.1 Search for the 511 keV from the star-forming regions
Figure 7: Same as Fig. 3 but with a reduced maximum intensity of $1\times 10^{-3}$ ph cm-2 s-1 sr-1 to enhance the visibility of faint structures.
Figure 8: (Top) Same as Fig. 7 but with annotations showing the regions used for flux measurements from individual sources (see Table 4). (Bottom) Same as the top panel but overlaid with contours of the sensitivity map shown in Fig. 2.
Here, we investigate the 511 keV emission from individual regions, particularly the Cygnus region and the Scorpius–Centaurus OB association (see Fig. 8). Including other star-forming regions, we evaluate the fluxes and their chance probabilities using the bootstrap samples with only the background.
@@
The Upper Scorpius region ($l=340^{\circ}$ to $360^{\circ}$, $b=25^{\circ}$ to $40^{\circ}$) is a part of the Scorpius–Centaurus association, the nearest OB association to the Sun. It shows a flux of $(1.4\pm_{\mathrm{stat}}0.8\pm_{\mathrm{sys}}0.5)\times 10^{-5}$ ph cm-2 s-1 and a chance probability of $8.0\times 10^{-2}$. This region corresponds to the chimney-like structure discussed in the previous subsection. Interestingly, 26Al maps measured by COMPTEL and INTEGRAL/SPI also show an enhanced emission from this region (Plüschke et al. 2001; Diehl et al. 2010). These findings suggest a possible spatial correlation between 26Al and positron annihilation. Considering that 26Al also emits a positron via its $\beta^{+}$ decay process, this correlation might shed light on the asymmetric vertical structure and the origin of the 511 keV emissions in this region with further confirmation. It is discussed more in Sect. 5, also regarding the “OSSE 511 keV fountain.”
We also investigated the 511 keV emission from other prominent star-forming regions, the Vela, Orion-Eridanus, Perseus, and Orion regions. The Vela region ($l=260^{\circ}$ to $290^{\circ}$, $b=-15^{\circ}$ to $15^{\circ}$) shows a large chance probability of 19% with an upper limit of $8.3\times 10^{-5}$ ph cm-2 s-1. The Orion region ($l=190^{\circ}$ to $220^{\circ}$, $b=-50^{\circ}$ to $10^{\circ}$) similarly shows no significant excess when considered alone. When extending the region to include the Perseus OB association, the chance probability is somewhat improved to 10.1%. It should be noted that there could be possible contamination from the statistical fluctuation of the nearby bright source, the Crab. When we extracted the bright region ($l=170^{\circ}$ to $200^{\circ}$, $b=-30^{\circ}$ to $0^{\circ}$), its flux was derived to be $6.5\pm_{\mathrm{stat}}2.4\pm_{\mathrm{sys}}1.9$ ph cm-2 s-1 with an improved chance probability of $1.2\times 10^{-2}$. A faint signal may be seen at $l=120^{\circ}$ close to the Galactic plane in Fig. 8, but its chance probability is about 60%, and we interpret it as a fluctuation in the reconstructed image.